# Structured Generative Modeling with The Thermodynamic Kolmogorov-Arnold Model

### v3

@@ Abstract
Learning an energy-based model (EBM) in the latent space of a top-down generative model offers an expressive and interpretable framework for text and image generation. However, it remains unclear how this interpretability can be used to guide model design, improve generative quality, and reduce training time. Moreover, the reliance on Langevin Monte Carlo (LMC) sampling presents challenges in efficiency and sampling multimodal latent distributions. In this work, we propose a novel adaptation of the Kolmogorov–Arnold representation theorem for generative modeling and introduce the Thermodynamic Kolmogorov–Arnold Model (T-KAM) to take advantage of structural and inductive biases. By constraining the prior to univariate relationships, T-KAM enables fast and exact inference via the inverse transform method. With the low dimensionality of the latent space and suitable inductive biases encoded, we demonstrate that importance sampling becomes a viable, unbiased, and highly efficient training strategy. We also introduce a training criterion using population-based LMC, which decomposes posterior sampling into a sequence of annealed distributions to improve multimodal exploration. T-KAM elegantly balances common trade-offs in generative modeling, offering fast inference, high sample quality, and stable training, while being naturally suited to upcoming Zettascale Computing Co. hardware and extendable to other high-impact research directions in generative intelligence.
@@ 1 Introduction
Firstly, MCMC is inherently iterative, and algorithms such as ULA require repeated gradient evaluations through the model, introducing significant computational overhead during both training and inference. Secondly, ULA is best suited for sampling from smooth, unimodal distributions and tends to struggle when exploring the multimodal latent distributions that are typical in latent variable generative modeling. Moreover, separate ULA samplers are required for prior and posterior sampling, which is problematic given that ULA is sensitive to hyperparameter tuning: step sizes that are too small lead to poor mixing and convergence, while larger step sizes can accumulate discretization bias due to its use of the Euler–Maruyama method to evolve the underlying Langevin equation.
@@
1.1 Inductive biases and importance sampling
To address these limitations, we reimagine the EBM prior framework into a novel univariate or element-wise model, drawing inspiration from Kolmogorov–Arnold Networks (KANs) (Liu et al., 2024) and deep spline networks Bohra et al. (2020), which replace the sigmoidal activations used in multilayer perceptrons Murtagh (1991) with spline interpolations. By representing the latent space with univariate functions and distributions, we improve interpretability for domain experts who may not be deeply familiar with computational statistics and machine learning.
This allows the EBM prior to be aligned with the posterior before training begins. We show that improved prior–posterior alignment enables efficient training with importance sampling (IS) (Goertzel, 1949), a technique for estimating posterior expectations that is otherwise avoided in the literature due to its degeneracy under severe prior–posterior mismatch and the curse of dimensionality, (Dupuis & Wang, 2004). Moreover, the trained prior can be visualized or recovered, offering new opportunities for transfer learning and scientific discovery.
This is tested in Sec. 3.1. Multiple KAN bases and reference priors are used to verify the effect of inductive bias. However, Gaussian Radial Basis Functions (RBF KANs) (Li, 2024), were used instead of the cubic B-spline bases proposed by Liu et al. (2024), as RBFs are more efficient for GPU implementation, and Poggio & Girosi (1990) previously demonstrated that they provide an optimal solution to the regularized approximation/interpolation problem.
To derive our new model, which we call the Thermodynamic Kolmogorov-Arnold Model (T-KAM), we first introduce a novel adaptation of the Kolmogorov–Arnold Representation Theorem (KART) (Givental et al., 2009) by reinterpreting its inner functions as Markov kernels between probability spaces. Given its finite and definitive nature, KART can provide structural bias and reduce the reliance of architecture design on heuristics and automated tuning. This reinterpretation allows us to frame inverse transform sampling (ITS) (Devroye, 1986) as a measurable transformation applied to uniform noise, thus preserving the deterministic structure of KART while admitting a probabilistic interpretation that aligns with MLE.
To address the potential inflexibility of strict adherence to KART and the limitations of using independent energy-based priors, we introduce a scalable strategy based on mixture distributions. Beyond introducing dependencies between latent dimensions to capture inter-dimensional relationships, this also improves sampling efficiency by enabling component-wise sampling.
@@ 1.2 Population-based Langevin sampling and Thermodynamic training
However, while ITS enables fast inference, posterior sampling during training must still be conducted using ULA when the IS estimator is insufficient. Prior work addressing poor MCMC convergence and multimodal exploration in EBMs has typically relied on diffusion models, such as the hierarchical EBM by Cui & Han (2024) and the diffusion-assisted EBM by Zhang et al. (2023). To maintain prior interpretability and fast inference post-training, we instead introduce posterior annealing as a novel and alternative strategy. This involves decomposing the posterior into a sequence of power posteriors, presented by Friel & Pettitt (2008), and sampling with population-based MCMC methods, specifically Parallel Tempering (PT) (Swendsen & Wang, 1986) (Swendsen & Wang, 1986; Marinari & Parisi, 1992; Hukushima & Nemoto, 1996).
@@
Although this may reduce expressivity compared to hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training. Our training criterion is then derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al. (2019) to ensure the impact of annealing on learning dynamics and parameter updates.
@@
While ULA is computationally efficient and enables more flexible chain dynamics, it does not preserve detailed balance or reversibility. This can be recovered by incorporating Metropolis-Hastings (MH) adjustments (Metropolis & Ulam, 1949) via the Metropolis-adjusted Langevin Algorithm (MALA) (Rossky et al., 1978a). The codebase in Sec. A.1 optionally supports MALA with locally adaptive step-size tuning, (autoMALA) (Biron-Lattes et al., 2024), to assist when a single step size fails to accommodate the varying geometries of power posteriors, and when the MH criterion deteriorates under the curse of dimensionality, (Hairer et al., 2014).
@@
MALA and autoMALA are discussed under Secs. A.6.3 and A.6.4. However, ULA remains the preferred method in our work, motivated by its demonstrated efficiency and effectiveness by Pang et al. (2020), and further supported by the findings of Nijkamp et al. (2019b; a), with regards to how short-run, biased MCMC methods can still be highly effective for sampling from EBMs.
@@
These scaling strategies are compared in Sec. 3.2. Despite our optimizations, the element-wise and localized nature of KAN functions complicate their parallelization on GPUs, which are best suited to matrix arithmetic. As a result, forward passes through T-KAM and the gradient evaluations required for ULA were found to be slow, necessitating reduced experimental complexity.
We ultimately found that our mixture prior model outperformed a deep KAN prior architecture closely aligned with the model used in the study by Pang et al. (2020), while also offering greater computational efficiency. In contrast, the use of annealing could not be justified due to its significant computational overhead and limited gains in quality.
@@
However, this assessment may change as a greater number of power posteriors and more sophisticated posterior sampling algorithms become feasible with the emerging hardware advances proposed by Zetta (2024). Their architecture enables fine-grained control over data flow and parallelized element-wise nonlinearity which is naturally well suited to supporting T-KAM at scale, while improving the potential of univariate priors with its function fusion features.
@@ 2 The Thermodynamic Kolmogorov-Arnold Model
This section introduces the model and its associated sampling procedures. We begin by presenting a novel interpretation of the Kolmogorov–Arnold Representation Theorem (KART) to define the model’s baseline architecture. Our key practical contribution is the introduction of a univariate energy-based prior, detailed in Sec. 2.1. The top-down generator can be instantiated in various forms, including Kolmogorov–Arnold Networks (KANs), which enables strict adherence to KART.
@@
The Kolmogorov–Arnold Representation theorem (KART) (Givental et al., 2009) forms the basis of T-KAM. A structured proof of the theorem has been provided by Dzhenzher & Skopenkov (2022).
@@
Eq. 1 can provide structural bias and guide the design of T-KAM’s architecture. To ensure adherence to it, any transformations must be derived to be continuous, measurable, and deterministic. For readers less familiar with the technical background, the two most practically significant equations of Sec.. 2.1 are Eqs. 46.
@@ 2.1 Univariate energy-based prior
is measurable. Here, $\delta_{\psi_{q,p}(u_{p})}$ denotes the Dirac measure centred at $\psi_{q,p}(u_{p})$, ensuring the kernel is deterministic: i.e., once $u_{p}$ is sampled, the outcome in $\mathcal{Z}$ is fixed. This measurability requirement guarantees that the pushforward from $([0,1],\mathcal{B}([0,1]),\mu)$ to $(\mathcal{Z},\mathcal{B}(\mathcal{Z}),\pi)$ is well defined and that no additional stochasticity is introduced beyond the uniform source variable, $u_{p}$. Here, $\mathcal{B}(\mathcal{Z})$ is the Borel $\sigma$-algebra on the latent space, $\mathcal{Z}$, and $\pi$ is the latent measure.
@@
In practice, $f_{q,p}$ can be parameterized in various ways, made possible by the recent development of Kolmogorov–Arnold Networks (KANs) (Liu et al., 2024). For instance, smoothing splines, such as those used by KANs or deep spline networks, (Bohra et al., 2020). Splines are typically defined on bounded domains, allowing the support of $\pi_{q,p}$ to be explicitly controlled by the user.
The partition function $Z_{q,p}$ is often intractable in energy-based models. However, in this univariate setting, where the energy function is smooth and defined on a bounded domain $\left[z_{\text{min}},z_{\text{max}}\right]$, it can be accurately approximated using Gauss–Kronrod quadrature, (Laurie, 1997):
@@ 2.3.1 Generator
Thus, a smooth transition between the latent space and data space can be achieved by progressively doubling layer widths, eliminating the need for arbitrary choices while maintaining relaxed compliance with KART. Strict adherence to KART can be achieved for hidden layer $h_{l_{k},l_{k-1}}$ by applying either a sigmoid or affine transformation after layer $h_{l_{k-2},l_{k-3}}$ to shift its range back to $0,1]^{L_{k-1}}$. However, using sigmoid may lead to vanishing gradients and we contend that function domains are less critical to KART than its structural properties and should not be prioritized over maintaining gradient flow and avoiding saturation.
@@ 2.3.2 Mixture prior
To encourage balanced weighting and reduce model complexity, L1 regularization can be applied to $\alpha_{q,p}$ and small values can be pruned. The mixture proportions are learned by adding $\log\alpha_{q,p}$ to the gradients of the prior in Sec. 2.4.1. Component selection follows a categorical distribution over $p=1,\dots,n_{z}$, determined by $\alpha_{q,p}$. Let $u_{\text{mix}}\sim\mathcal{U}(u_{\text{mix}};\;0,1)$. Guided by Devroye (2006) regarding how ITS applied to discrete variables, the selected component $p^{*}$ is given by:
@@ 2.3.3 Deep KAN prior
If required, we may further deepen the energy functions $f_{q,p}$ using the KAN approach. In fact, each $f_{q,p}$ can itself be a realization of Eq. 1 when its domain is restricted to $[0,1]$. This extension introduces new challenges:
@@
where $\bm{z}\in\mathcal{Z}\subset\mathbb{R}^{n_{z}}$, and is constrained to $0,1]^{n_{z}}$ when strict adherence to KART is required. Sampling from this prior can be carried out using the Unadjusted Langevin Algorithm (ULA), as described in Sec. 2.5.2.
@@ 2.4.4 Posterior Expectations with Importance Sampling
    Otherwise, the weights are taken without alteration, (i.e. the next step is skipped). When the weights are uniformly spread the ESS evaluates to $N_{s}$. High-variance weights with a low ESS indicate that many samples contribute minimally, resulting in inefficient sampling. Resampling focuses learning on samples that are more representative of the posterior distribution, thereby maximizing the usefulness of each parameter update.
@@
This estimator is used for computing the log-marginal likelihood and its gradient in Eqs. 15 and 16. However, unlike ULA (discussed in Sec. 2.5.2 and demonstrated in Sec. 3.2), IS fails to support posterior exploration when the prior is poorly aligned, making it unsuitable for general-purpose tasks beyond those similar to the datasets presented in Secs. 3.1.

### v4

@@ Abstract
Learning an energy-based model (EBM) in the latent space of a top-down generative model offers a versatile framework for generation across multiple data modalities. However, it remains unclear how its interpretability can be used to guide model design, improve generative quality, and reduce training time. Moreover, the reliance on Langevin Monte Carlo (LMC) sampling presents challenges in efficiency and sampling multimodal latent distributions. In this work, we propose a novel adaptation of the Kolmogorov-Arnold representation theorem for generative modeling and introduce the Thermodynamic Kolmogorov-Arnold Model (T-KAM) to take advantage of structural and inductive biases. By constraining the prior to univariate relationships, T-KAM enables fast and exact inference via the inverse transform method. With the low dimensionality of the latent space and suitable inductive biases encoded, we demonstrate that importance sampling (IS) becomes a viable, unbiased, and highly efficient posterior sampler. For situations where IS fails, we investigate a novel strategy using population-based LMC, which decomposes posterior sampling into a sequence of annealed distributions to improve multimodal sampling. T-KAM elegantly balances common trade-offs in generative modeling, offering fast inference, interpretability, and stable training, while being naturally suited to upcoming Zettascale Computing Corp. hardware.
@@ 1 Introduction
Firstly, MCMC is inherently iterative, and algorithms such as ULA require repeated gradient passes through the model, which introduces significant computational overhead during both training and inference. Secondly, ULA is best suited for sampling from smooth, unimodal distributions and can struggle to explore the multimodal latent distributions that are typical in latent variable generative modeling. Moreover, the framework presented by Pang et al. (2020) requires separate ULA samplers for the prior and posterior. This introduces additional experimental complexity since ULA is sensitive to hyperparameter tuning: step sizes that are too small lead to poor mixing and convergence, while larger step sizes can accumulate discretization bias due to its use of the Euler-Maruyama method (Wilkie, 2004) to evolve the underlying Langevin equation.
@@
1.1 Inductive biases and fast sampling
To address these limitations, we reimagine the EBM prior into a novel univariate or element-wise model, drawing inspiration from Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024) and deep spline networks Bohra et al. (2020), which replace the sigmoidal activations used in multilayer perceptrons Murtagh (1991) with spline interpolations. By representing the latent space with univariate functions and distributions, we enable fast and exact prior sampling with the inverse transform method (Devroye, 1986) and improve interpretability for domain experts who may not be deeply familiar with computational statistics and machine learning. This allows the EBM prior to be aligned with the posterior before training begins.
In Sec. 3.1, we show that adequately chosen priors can enable efficient training with importance sampling (IS) (Goertzel, 1949). This is a technique for estimating posterior expectations that is otherwise avoided in the literature due to its high variance under severe prior-posterior mismatch and the curse of dimensionality, (Dupuis & Wang, 2004), which can be avoided in our model due to the low dimensionality of the latent space. Moreover, we show that the trained priors are easy to visualize and recover, offering new opportunities for scientific discovery and knowledge transfer to models with different latent structures.
Multiple KAN bases and reference priors were used to verify the effect of inductive bias on training with IS. However, Gaussian Radial Basis Functions (RBF KANs) (Li, 2024), were adopted instead of the cubic B-spline bases proposed by Liu et al. (2024), as RBFs are more efficient for GPU implementation, and Poggio & Girosi (1990) previously demonstrated that they provide an optimal solution to the regularized approximation/interpolation problem.
To derive our new model, which we call the Thermodynamic Kolmogorov-Arnold Model (T-KAM), we first introduce a novel adaptation of the Kolmogorov-Arnold Representation Theorem (KART) (Givental et al., 2009) by reinterpreting its inner functions as Markov kernels between probability spaces. Given its finite and definitive nature, KART can provide structural bias and reduce the reliance of architecture design on intuition and automated tuning.
This reinterpretation allows us to frame inverse transform sampling (ITS) as a measurable transformation applied to uniform noise, thus preserving the deterministic structure of KART while admitting a probabilistic interpretation that aligns with MLE. To address the potential inflexibility of strict adherence to KART and the limitations of using independent energy-based priors, we introduce a scalable strategy based on mixture distributions. Beyond introducing dependencies between latent dimensions to capture inter-dimensional relationships, this also improves sampling efficiency by enabling component-wise sampling.
@@ 1.2 Population-based Langevin sampling and Thermodynamic training
However, while ITS enables fast inference, posterior sampling must still be conducted using ULA when the IS estimator is insufficient. Prior work addressing poor MCMC convergence and multimodal exploration in EBMs has typically relied on diffusion models, such as the hierarchical EBM by Cui & Han (2024) and the diffusion-assisted EBM by Zhang et al. (2023). To maintain prior interpretability and fast inference post-training, we instead introduce posterior annealing as a novel and alternative strategy. This involves decomposing the posterior into a sequence of power posteriors, presented by Friel & Pettitt (2008), and recasting ULA with population-based MCMC methods (Swendsen & Wang, 1986; Marinari & Parisi, 1992; Hukushima & Nemoto, 1996).
@@
The resulting training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al. (2019) to ensure the impact of annealing on learning dynamics and parameter updates. Although this reduces expressivity compared to hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, annealing is embarrassingly parallel.
@@
While ULA is computationally efficient and enables more flexible chain dynamics, it does not preserve detailed balance or reversibility. This can be recovered by incorporating Metropolis-Hastings (MH) adjustments (Metropolis & Ulam, 1949) via the Metropolis-adjusted Langevin Algorithm (MALA) (Rossky et al., 1978a). The code in Sec. A.1 optionally supports MALA with locally adaptive step-size tuning, (autoMALA) (Biron-Lattes et al., 2024), to assist when a single step size fails to accommodate the varying geometries of power posteriors, and when the MH criterion deteriorates under the curse of dimensionality, (Hairer et al., 2014). This is discussed in Secs. A.6.3 and A.6.4.
@@
However, ULA remains the preferred method in our work, motivated by its efficiency and effectiveness, demonstrated by Pang et al. (2020) and further supported by the findings of Nijkamp et al. (2019b; a), with regards to how short-run, biased MCMC methods can still be highly effective for sampling from EBMs.
@@
MLE and SE are compared in Sec. 3.2. Despite our optimizations, the element-wise and localized nature of KANs complicate their parallelization on GPUs, which are best suited to matrix arithmetic. As a result, forward passes through T-KAM and the gradient evaluations required for ULA were found to be slow, necessitating reduced experimental complexity.
@@
The experiments in Sec. 3.2 demonstrate that our mixture prior model is able to generate coherent, (although low-quality), samples after a few training epochs on RGB images, while still supporting efficient, exact sampling via component-wise ITS. In contrast, annealing incurred substantial computational overhead, but produced little to no improvements in sample quality, rendering its use difficult to justify. However, this result may change as a larger number of power posteriors, longer training durations, and more advanced posterior sampling algorithms become practical with the hardware advances proposed by Zetta (2024). Their architecture enables fine-grained control over data flow and parallelized element-wise nonlinearity which is naturally well suited to supporting T-KAM at scale, while improving the potential of univariate priors with its function fusion features.
@@ 2 The Thermodynamic Kolmogorov-Arnold Model
This section introduces the model and its associated sampling procedures. We begin by presenting a novel interpretation of the Kolmogorov-Arnold Representation Theorem (KART) to define the model’s baseline architecture. Our key practical contribution is the introduction of a univariate energy-based prior, detailed in Sec. 2.1. The top-down generator can be instantiated in various forms, including Kolmogorov-Arnold Networks (KANs), which enables strict adherence to KART.
@@
The Kolmogorov-Arnold Representation theorem (KART) (Givental et al., 2009) forms the basis of T-KAM. A structured proof of the theorem has been provided by Dzhenzher & Skopenkov (2022).
@@
Eq. 1 can provide structural bias and guide the design of T-KAM’s architecture. To ensure adherence to it, any transformations must be derived to be continuous, measurable, and deterministic. For readers less familiar with the technical background, the two most practically significant equations of Sec. 2.1 are Eqs. 46.
@@ 2.1 Univariate energy-based prior
is measurable. Here, $\delta_{\psi_{q,p}(u_{p})}$ denotes the Dirac measure centred at $\psi_{q,p}(u_{p})$, ensuring the kernel is deterministic: i.e., once $u_{p}$ is sampled, the outcome in $\mathcal{Z}$ is fixed. This guarantees that the pushforward from $([0,1],\mathcal{B}([0,1]),\mu)$ to $(\mathcal{Z},\mathcal{B}(\mathcal{Z}),\pi)$ is well defined and that no additional stochasticity is introduced beyond the uniform source variable, $u_{p}$. Here, $\mathcal{B}(\mathcal{Z})$ is the Borel $\sigma$-algebra on the latent space, $\mathcal{Z}$, and $\pi$ is the latent measure.
@@
In practice, $f_{q,p}$ can be parameterized in various ways, made possible by the recent development of Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024). For instance, smoothing splines, such as those used by KANs or deep spline networks, (Bohra et al., 2020). Splines are typically defined on bounded domains, allowing the support of $\pi_{q,p}$ to be explicitly controlled by the user.
The partition function $Z_{q,p}$ is often intractable in energy-based models. However, in this univariate setting, where the energy function is smooth and defined on a bounded domain $\left[z_{\text{min}},z_{\text{max}}\right]$, it can be accurately approximated using Gauss-Kronrod quadrature, (Laurie, 1997):
@@ 2.3.1 Generator
Thus, a smooth transition between the latent space and data space can be achieved by progressively doubling layer widths, eliminating the need for arbitrary choices while maintaining relaxed compliance with KART. Strict adherence to KART can be achieved for hidden layer $h_{l_{k},l_{k-1}}$ by applying either a sigmoid or affine transformation after layer $h_{l_{k-2},l_{k-3}}$ to shift its range back to $0,1]^{L_{k-1}}$. However, using sigmoid may lead to vanishing gradients and we suggest that function domains are less critical to KART than its structural properties and should not be prioritized over maintaining gradient flow and avoiding saturation.
@@ 2.3.2 Mixture prior
The mixture proportions are learned by replacing the log-prior in Eq. 18 with the logarithm of Eq.12. To encourage balanced weighting and reduce model complexity, L1 regularization can be applied to $\alpha_{q,p}$ and small values can be pruned.
Component selection follows a categorical distribution over $p=1,\dots,n_{z}$, determined by $\alpha_{q,p}$. Let $u_{\text{mix}}\sim\mathcal{U}(u_{\text{mix}};\;0,1)$. Guided by Devroye (2006) regarding how ITS applied to discrete variables, the selected component $p^{*}$ is given by:
@@ 2.3.3 Deep KAN prior
If required, the energy functions $f_{q,p}$ can be deepened using the KAN approach. In fact, each $f_{q,p}$ can itself be a realization of Eq. 1 when its domain is restricted to $[0,1]$. This extension introduces new challenges:
@@
where $\bm{z}\in\mathcal{Z}\subset\mathbb{R}^{n_{z}}$, and is constrained to $0,1]^{n_{z}}$ when strict adherence to KART is required. Sampling from this prior can be carried out using the Unadjusted Langevin Algorithm (ULA), as described in Sec. 2.5.2. This is not investigated in the experiments, since it is simply equivalent to the model of Pang et al. (2020), reformulated as a KAN.
@@ 2.4.2 Log-likelihood
The code in Sec. A.1 also provides an option to augment the first likelihood model with an additional perceptual loss term (Johnson et al., 2016) using features extracted from the pretrained VGG-16 CNN (Simonyan & Zisserman, 2015). This could have yielded higher-fidelity images for Sec. 3.2, but it was not adopted in this study due to the computational overhead of Langevin sampling.
@@ 2.4.4 Posterior Expectations with Importance Sampling
    Otherwise, the weights are taken without alteration, (i.e. the next step is skipped). When the weights are uniformly spread the ESS evaluates to $N_{s}$. High-variance weights with a low ESS indicate that many samples contribute minimally, resulting in inefficient sampling. Resampling focuses learning on samples that are more representative of the posterior distribution, which maximizes the usefulness of each parameter update.
@@
This is used to estimate the log-marginal likelihood and its gradient in Eqs. 15 and 16. However, unlike ULA (discussed in Sec. 2.5.2 and demonstrated in Sec. 3.2), IS fails to support posterior exploration when the prior is poorly aligned, making it unsuitable for general-purpose tasks beyond those similar to the datasets presented in Secs. 3.1.